# Examples of robotic art Encountering Robotic Art: The Social, Material, and Temporal Processes of Creation with Machines

### v1

@@ Abstract.
Robots extend beyond the tools of productivity; they also contribute to creative activities. Although typically defined as utility-driven technologies designed for productive or social settings, the role of robots in creative settings remains underexplored. This paper examines how robots participate in artistic creation. Through semi-structured interviews with robotic artists, we analyze the impact of robots on artistic processes and outcomes. We identify the critical roles of social interaction, material properties, and temporal dynamics in facilitating creativity. Our findings reveal that creativity emerges from the co-constitution of artists, robots, and audiences within spatial-temporal dimensions. Based on these insights, we propose several implications for socially informed, material-attentive, and process-oriented approaches to creation with computing systems. These approaches can inform the domains of HCI, including media and art creation, craft, digital fabrication, and tangible computing.
@@ 1. Introduction
The term “robot,” originally signifying ‘forced labor,’ has long been associated with labor and work. Robots have demonstrated their utility in various automated productive and social contexts, where the primary goals are improving productivity, safety, and fostering social interactions with humans (Simões et al., 2022; Weidemann and Rußwinkel, 2021; Honig and Oron-Gilad, 2018). However, an increasing number of cases feature the use of robots in creative settings. Unlike productive contexts, where the focus is on efficiency and task completion (Arents and Greitans, 2022), or social contexts, where communication and trust are prioritized (Nam and Lyons, 2020; Saunderson and Nejat, 2019), creative environments prioritize artistic innovation and expression (Hsueh et al., 2024). This shift fundamentally alters the dynamics of human-robot interaction, redefining the roles and expectations for both humans and robots.
@@
For instance, robots’ social behaviors are leveraged to support the generation and expression of creative ideas (Hu et al., 2021; Sandoval et al., 2022; Alves-Oliveira et al., 2020), while programmable robotic movements and trajectories are employed to inspire artistic activities such as sketching (Lin et al., 2020). These studies often engage participants from creative fields who possess limited prior experience with robotics, and are typically conducted in short-term, experimental settings. Consequently, the findings from these studies remain constrained since much can be learned from professional practitioners’ experiences to inform system design such as digital fabrication (Hirsch et al., 2023). There is a notable gap in research examining artists with long-term, active, and practical experience in integrating robotic systems into their creative processes. As a result, the deeper insights into how robots facilitate and shape creative processes, beyond simply augmenting human creativity, remain underexplored. In this study, we aim to better understand the impacts of robots on creative processes and outcomes.
@@
As early as Leonardo da Vinci’s 16th century “Automaton,” artists have explored the creative affordances of robotic systems (Shanken, 2002; Pagliarini and Hautop Lund, 2009; Jeon, 2017). The artistic creation process typically encompasses various stages, including the exploration of materials and techniques, ongoing experimentation and iteration, and the continual refinement of the artists’ insights into their creative subjects (Lewis et al., 2023; Sturdee et al., 2022). Therefore, investigating the artistic process involving robots offers a valuable opportunity to gain deeper insights into their creative potential. Robotic art, in particular, provides a compelling context for this exploration.
@@
Robotic art refers to artworks that utilize robotic or automated technologies to create artistic experiences and tangible artifacts. One example is robotic installation art, where robots are programmed to follow specific rules that embody the artist’s expression (Fig. 1 (a)). Another example is responsive art, in which robots react to their environment, with behaviors that change over time or in response to spectators (Fig. 1 (b)). Additionally, there are robotic creators, which possess a degree of autonomy, allowing them to collaborate with human artists and produce works that extend beyond mere replication of human-created art (Fig. 1 (c and d)). As such, robotic art offers a rich case for exploring human-machine interactions in creative contexts. Gaining a deeper understanding of how robots facilitate artistic expression can provide valuable insights for designing computing systems to support creative activities (Gomez Cubero et al., 2021).
@@
We draw on semi-structured, in-depth interviews with renowned professional robotic artists to investigate the interactive patterns associated with the use of robots in artistic practice. Specifically, our goal is to understand how artistic exploration of robotic systems challenges conventional assumptions about the functions of robots, such as their roles in automating repetitive tasks or serving human needs. We also explore the implications of robots in artwork and examine how creativity emerges within robotic art. To address these interrelated issues, our study focuses on the practice of robotic art, posing the research question: How do robotic artists utilize robots in their artistic practice? We approach this inquiry through the perspectives and experiences of robotic artists, who creatively design, modify, and repurpose robotic systems for artistic expression and exploration.
@@
Our findings highlight the social, material, and temporal dimensions of artists’ practices that shape their creativity and artistic outcomes. The creation of robotic art is largely a social process, as artists receive both explicit and implicit feedback through the audience’s reactions and reception of their work. Simultaneously, the embodiment and malfunctions inherent to robotic systems drive artistic experimentation. Moreover, the temporal processes of creation and exhibition, beyond just the final product, further enhance the creative value. Our empirical analysis demonstrates how creativity emerges through the interplay of social, material, and temporal interactions among artists, robots, audiences, and the environment.
@@
We make two main contributions to HCI in this study. First, we elucidate the interactive mechanisms among key actors—human creators, machines, audiences, and environments—within the practice of robotic art, an area that remains underexplored in HCI. Our findings reveal the significance of sociality (e.g., interactions between artists and audiences), materiality (e.g., the embodiment and malfunctions of robots), and temporality (e.g., the processes of creation and exhibition) in shaping creative values. We propose that these three facets are central to the creative process and facilitate the emergence of creativity in robotic art. Second, drawing from the findings, we offer implications for socially informed, material-attentive, and process-oriented creation with computing systems. We suggest leveraging these three aspects to enhance creativity and the creative experience. Specifically, we discuss the value of incorporating implicit audience feedback, designing for technical malfunctions, and focusing on post-creation process to foster alternative creative experiences with machines (Alter, 2010; Juarez, 2022).
@@ 2. Related Work
As we study practices of artists making robotic art with the intent to inform ways of facilitating creativity, we review literature about robotic art as a case of creative practice to probe us to understand the factors that contribute to perceived creativity in the artwork. To justify our contributions as learning from the art to inform design, we review artistic practices in HCI and their unique benefits to the field.
@@ 2.1. Robots in Creative Contexts
Introducing robots into creative domains provides a new research medium through robots’ embodied materiality and interactions (Nitsche and Weisling, 2019). Robots applied in creative tasks have been treated not only as executors following a prescribed set of predetermined rules to execute commands but also as social actors that can respond to dynamic human behaviors and environments (Nam et al., 2023). For example, robot collaborative drawing highlights the robot’s role as a tool, to mediate and engage the spatial arrangement, physical activities, and social interactions of people in the co-creation process, through robotic motions and creations (Hoggenmueller and Hespanhol, 2020). The collaboration between humans and robots in choreography provides a lens to explore how human movement can be reinterpreted and fused by robots, thus inspiring more creative expression (Saviano et al., 2024; Hsueh et al., 2019b). In this context, robots function not merely as tools but as partners that contribute to the improvisations (Rogel, 2022). Thörn et al. examined the potential of robots as dance partners (Thörn et al., 2020), emphasizing the importance of synchronization of robotic movement and emotional expression in creating a compelling dance experience (Nam and Nitsche, 2014). The participation of social robots in storytelling underscored the potential of robots to enhance creative processes through their presence, body language, facial expressions, and vocal modulations (Jeon, 2017). Antunes et al. (2022) demonstrated how robots can act as co-narrators or audience members in storytelling sessions, thereby enriching the narrative experience and fostering children’s creativity through facilitating multisensory feedback and emotional behaviors.
@@
These examples illuminate the multifaceted roles that robots can play in creative processes, from facilitating creative processes as tools and partners to acting as resources (Chung et al., 2021). To fully grasp the potential of robots in the creative domain, it is essential to explore how their material features and interactive capabilities shape and support creative expression. Beyond the roles that robots play in creative processes, we are also interested in what is unique about robots (e.g., material features and interaction processes) that accounts for the effort to keep advancing and applying them in creative contexts in a more integrated way. The artistic appropriation of robots may provide part of the answer. Notably, artists engage proactively with robots when creating artwork, exploiting the artistic potential of robots, rather than passively accommodating the robots’ inherent limitations. They break down robots, redefine them, and repurpose them as instruments for artistic goals. We propose that robots’ specific manifested qualities hold the potential to stimulate creativity and drive such interest and engagement by artists in their practice. Thus, we aim to better answer the above questions by examining the practices and characteristics of artists’ voluntary utilization of robots in achieving artistic ends. Answering these questions can contribute to knowledge about the unique capabilities and roles of robots in creative contexts.
@@ 2.2. Robotic Art
Robotic art, a term that encompasses the intersection of machine intelligence and creativity, involves the use of robots as both mediums and subjects of artistic expression (Penny, 2016; Kac, 1997). Robotic art covers a group of practices that is usually related to computer-automated cultural artifacts (CACA) (Penny, 2013), although some authors refer to the broader field of mechanically driven kinetic art in an attempt to frame it within media art practices (Hoetzlein, 2009). From a general perspective, a collection of works with a certain level of machine autonomy is often addressed as robotic art (Kac, 2001), belonging to a broader category of machinic art, which includes ‘telerobotic art’ (Reichle et al., 2009), ‘postbiological life forms’ (Reichle et al., 2009), ‘robotic installation art’ (Shi, 2023), and so on. Artworks featuring robots not only create more engaging experiences, but also challenge the boundaries of intelligent agency and human creativity. For instance, robotic artists prompt us to question whether a robot that creates art can be considered an artist in its own right (Shi, 2023; Mikalonytundefined and Kneer, 2022).
Robots have been adopted in artistic expression and creation due to their anthropomorphic qualities, creating a metaphor as social components and providing a degree of sociality (Jeon, 2017; Kahn et al., 2014). In this context, artists can leverage robots’ social behaviors to stimulate artistic expression from the creators. For instance, natural language dialogue can establish social connections with creative producers and guide creative imagination through questioning (Kahn et al., 2016; Hu et al., 2021). Artists often utilize robots’ social behaviors such as facial expressions, body postures, movements, and speech intonation to provide emotional communication during the creative theater performances, thereby encouraging the generation of creative human-human/robot interaction (Knight, 2011; Li et al., 2023). These characteristics distinguish robots from traditional artistic materials such as plaster or fabric, allowing artists to create artwork and express themselves more dynamically through interactions with living materials.
@@
Thanks to robots’ interactivity, robotic artworks often incorporate audiences into the creation process, allowing artists to harness more collective creativity and enabling greater possibilities in artistic expression (Guljajeva and Canet Sola, 2022; Mikalauskas et al., 2018). Examples include Tresset’s performance installations with Paul, a custom robot art and computer program for real-time portrait drawing (Tresset and Deussen, 2014). In these works, audiences not only observe a robot creating art, but are invited to witness the creative process of a technological ensemble at work, observing how human artists and robotic systems continuously influence each other (Chung, 2022; Gomez Cubero et al., 2021; Tresset and Deussen, 2014).
While these researches have demonstrated the potential of robots in artistic creation and their capacity to stimulate creativity, they discuss discretely how creativity is supported and unleashed under specific robot designs within individual robotic systems, revealing limited approaches toward creativity. There remains a need for research that explores multiple cases involving diverse manifestations of robots to deepen our understanding of how robots can contribute meaningfully to the creative process. Our study addresses this gap by analyzing the integrated creative practices of multiple robotic art practitioners across various mediums, such as drawing, dance, and installation. This approach allows us to uncover shared experiences and insights across different artistic forms, potentially revealing broader principles for how robots can effectively facilitate and shape creative practices.
@@ 2.3. Artistic Practices in HCI
The relationship between art and technology is becoming increasingly intertwined, with artistic practices being extensively integrated into the HCI community to study human interactions with technology and machines in specific contexts (Jeon et al., 2019). By adopting exploratory approaches, these practices challenge traditional usability study methods such as scale measurements. For instance, some studies involve participants in spatial interactive sound installations, using sound theory and micro-phenomenological interviews to analyze the interplay between sensory perception, bodily intent, and sound (Robson et al., 2024; Frid et al., 2019). Through these methods, researchers can capture the subtleties of how users perceive and engage with sound in immersive environments that are often overlooked by conventional measurement tools. Rajcic and McCormack (2020) introduced an emotion-responsive mirror that generates poetry in real-time, exploring the contrasts in emotional cognition, processing, and expression between machines and users. Similarly, Being the Machine by Devendorf and Ryokai (2015) invites artists to perform 3D printing manually in a machine-like manner, probing questions of agency and control in human-machine interactions. Furthermore, collaborations with choreographers have led to the creation of movement-sonification scarves as probes in dance performances, investigating dancers’ engagement with music and aiding HCI researchers in exploring how to design body-centric interaction technologies (Bang et al., 2023). The creation of artistic works, with their inherent ambiguity and potential for unforeseen human interactions, parallels the challenges in studying human interactions with emerging technologies (Coessens et al., 2009). Interactive art pieces and open exploratory spaces empower HCI researchers to expand their understanding of emergent human behaviors.
@@
However, the role of artistic practice in HCI research has been a subject of ongoing debate. Some researchers question integrating artistic knowledge into HCI (Frankjaer and Dalsgaard, 2020; Candy, 2011). However, examining the relationship between art and technology serves as both inspiration and a means to understand the interaction and the broader social world (Taylor et al., 2017), leading to novel design research methods such as cultural probes (Gorichanaz, 2020; Gaver et al., 1999), participatory/co-design (Clarke et al., 2016; DiSalvo et al., 2007; Soden et al., 2020), and arts-based ethnography (Kang et al., 2022). Some studies incorporate interdisciplinary teams in the creation and performance of artworks, using participant observation and interviews to explore the nonlinear collaborative practices across disciplines, highlighting the intersection of technological and aesthetic thinking (Kang et al., 2022, 2018). Research on self-situated performance practices, through interviews of artists, has expanded our understanding of new relationships and interactions between researchers and participants in HCI, allowing for ongoing refinement and enhancement of research practices (Taylor et al., 2017). In these cases, artistic practice is not merely an object of interpretation but shapes new exploratory processes, offering creators and participants opportunities for reflection and knowledge transformation.
@@
Besides, the artistic practices also bring diverse aesthetic experiences to the HCI community, helping researchers focus more on the sensory dimensions of interactions with technology. The concept of aesthetic computing, exploring how art and aesthetics are applied in different areas of computer science, introduces a sensuous aesthetic framework to the exploration of technology in HCI research (Fishwick, 2006; Jeon, 2017). Artistic practices bring embodiment and interaction to the forefront, placing human sensory experience, judgment, and contemplation at the core of aesthetic experiences (Taylor et al., 2017). For example, soma design focuses on elevating the sensory experiences of bodily activities to consciousness, utilizing these experiences as resources in the design process. Designers engage deeply with materials—whether physical, digital, cultural, or movement-based—through all their senses, cultivating their aesthetic sensibilities to create innovative designs (Hook, 2018; Höök et al., 2016; Luft et al., 2023). Consequently, various artistic practices, including choreography (Allen d’Ávila Silveira et al., 2022; Eriksson et al., 2019), sound art (Kilic Afsar et al., 2023), and media art (Branch et al., 2021), are increasingly incorporated to help researchers and users refocus on bodily sensations, prompting a reevaluation of our judgments about technology.
@@
Integrating art practices into HCI research has shown significant potential in broadening our understanding of user experiences and technology design. Our study explores how robotic art can help the HCI community deepen its understanding of the relationship between human creative behaviors and robots as exploratory materials. We aim to bridge the gap between artistic practices and technology designs for creativity by proposing a framework explicating how computing systems like robots serve as the tools and sites for creative endeavors and offering subsequent design implications for creative activities.

### v2

@@ Abstract.
Robots extend beyond the tools of productivity; they also contribute to creativity. While typically defined as utility-driven technologies designed for productive or social settings, the role of robots in creative settings remains underexplored. This paper examines how robots participate in artistic creation. Through semi-structured interviews with robotic artists, we analyze the impact of robots on artistic processes and outcomes. We identify the critical roles of social interaction, material properties, and temporal dynamics in facilitating creativity. Our findings reveal that creativity emerges from the co-constitution of artists, robots, and audiences within spatial-temporal dimensions. Based on these insights, we propose several implications for socially informed, material-attentive, and process-oriented approaches to creation with computing systems. These approaches can inform the domains of HCI, including media and art creation, craft, digital fabrication, and tangible computing.
@@ 1. Introduction
The term “robot,” originally signifying ‘forced labor,’ has long been associated with labor and work. Robots have demonstrated their utility in various automated productive and social contexts, where the primary goals are improving productivity, safety, and fostering social interactions with humans (Simões et al., 2022; Weidemann and Rußwinkel, 2021; Honig and Oron-Gilad, 2018). However, an increasing number of cases feature using of robots in creative settings. Unlike productive contexts, where the focus is on efficiency and task completion (Arents and Greitans, 2022), or social contexts, where communication and trust are prioritized (Nam and Lyons, 2020; Saunderson and Nejat, 2019), creative environments prioritize artistic innovation and expression (Hsueh et al., 2024). This shift fundamentally alters the dynamics of human-robot interaction, redefining the roles and expectations for both humans and robots.
@@
For instance, robots’ social behaviors are leveraged to support the generation and expression of creative ideas (Hu et al., 2021; Sandoval et al., 2022; Alves-Oliveira et al., 2020), and programmable robotic movements and trajectories are employed to inspire artistic activities such as sketching (Lin et al., 2020). These studies often engage participants from creative fields who possess limited prior experience with robotics, and are typically conducted in short-term, experimental settings. Consequently, the findings from these studies remain constrained since much can be learned from professional practitioners’ experiences to inform system design such as digital fabrication (Hirsch et al., 2023). There is a notable gap in research examining the long-term, active, and practical experience of integrating robotic systems into the creative processes. As a result, the deeper insights into how robots facilitate and shape creative processes, beyond simply augmenting human creativity, remain underexplored. In this study, we aim to better understand the impacts of robots on creative processes and outcomes.
@@
As early as Leonardo da Vinci’s 16th century “Automaton,” artists have explored the creative affordances of robotic systems (Shanken, 2002; Pagliarini and Hautop Lund, 2009; Jeon, 2017). The artistic creation process typically encompasses various stages, including the exploration of materials and techniques, ongoing experimentation and iteration, and the continual refinement of the artists’ insights into their creative subjects (Lewis et al., 2023; Sturdee et al., 2022). Therefore, investigating the artistic process involving robots offers an opportunity to gain deeper insights into robots’ creative potential. Robotic art, in particular, provides a compelling case for this exploration.
@@
We define robotic art as artworks that utilize robotic or automated machines to create artistic experiences and tangible artifacts. One example is robotic installation art, in which robots are programmed to follow specific rules that embody the artist’s expression (Figure 1 (a)). Another example is responsive art, in which robots react to their environment, with behaviors that change over time or in response to spectators (Figure 1 (b)). Additionally, there are robotic creators, which possess a degree of agency, allowing them to collaborate with human artists and produce works that extend beyond mere replication of human-created art (Figure 1 (c) and (d)). As such, robotic art becomes a rich case for exploring human-machine interactions in creative contexts. Gaining a deeper understanding of how robots facilitate artistic expression can provide insights for designing computing systems to support creative activities (Gomez Cubero et al., 2021).
@@
We draw on semi-structured, in-depth interviews with renowned professional robotic artists to investigate the use of robots in artistic practice. Specifically, our goal is to understand how artistic exploration of robotic systems challenges conventional assumptions about the functions of robots, such as their roles in automating repetitive tasks or serving human needs. We also explore the implications of robots in the artistic process and examine how creativity may emerge within robotic art. To address these interrelated inquiries, our study focuses on the practice of robotic art, posing the research question: How do robotic artists utilize robots in their artistic practice? We approach this inquiry through the perspectives and experiences of robotic artists, who creatively design, modify, and repurpose robotic systems for artistic expression and exploration.
@@
Our findings highlight the social, material, and temporal dimensions of artists’ practices that shape their creativity and artistic outcomes. The creation of robotic art is largely a social process, as artists receive both explicit and implicit feedback through the audience’s reactions and reception of their work. Simultaneously, the embodiment and malfunctions inherent to robotic systems drive artistic experimentation. The temporal processes of creation and exhibition, beyond just the final product, further enhance the creative value. Our empirical analysis presents how creativity emerges through the interplay of social, material, and temporal interactions among artists, robots, audiences, and the environment.
@@
We make two main contributions to HCI in this study. First, we elucidate the interactive mechanisms among key actors—human creators, machines, audiences, and environments—within the practice of robotic art, a topic that remains underexplored in HCI. Our findings reveal the significance of sociality (e.g., interactions between artists and audiences), materiality (e.g., the embodiment and malfunctions of robots), and temporality (e.g., the processes of creation and exhibition) in shaping creative values. We propose that these three facets are central to the creative process and facilitate the emergence of creativity in robotic art. Second, drawing from the findings, we offer implications for socially informed, material-attentive, and process-oriented creation with computing systems. We suggest leveraging these three aspects to enhance creativity and the creative experience. Specifically, we discuss the value of incorporating implicit audience feedback, designing with technical malfunctions, and focusing on the post-creation process to foster alternative creative experiences with machines (Alter, 2010; Juarez, 2022).
@@ 2. Related Work
As we study practices of artists making robotic art with the intent to inform ways of facilitating creativity, we review literature about robotic art as a case of creative practice to probe the factors that contribute to creativity in the artwork. To justify our contributions as learning from the art to inform design, we review artistic practices in HCI and their unique benefits to the field.
@@ 2.1. Robots in Creative Contexts
Introducing robots into creative domains provides a new research medium through robots’ embodied materiality and interactions (Nitsche and Weisling, 2019). Robots applied in creative tasks have been treated not only as executors following a set of predetermined rules and executing commands but also as social actors that can respond to dynamic human behaviors and environments (Nam et al., 2023). For example, robot collaborative drawing highlights robot’s role as a tool, to mediate and engage the spatial arrangement, physical activities, and social interactions of people in the co-creation process through robotic motions and creations (Hoggenmueller and Hespanhol, 2020). The collaboration between humans and robots in choreography provides a lens to explore how human movement can be reinterpreted and fused by robots, thus inspiring more creative expression (Saviano et al., 2024; Hsueh et al., 2019b). In this context, robots function not merely as tools but as partners that contribute to improvisation (Rogel, 2022). Thörn et al. (Thörn et al., 2020) examined the potential of robots as dance partners, emphasizing the importance of synchronization of robotic movement and emotional expression in creating a compelling dance experience (Nam and Nitsche, 2014). The participation of social robots in storytelling underscored the potential of robots to enhance creative processes through their presence, body language, facial expressions, and vocal modulations (Jeon, 2017). Antunes et al. (2022) demonstrated how robots can act as co-narrators or audience members in storytelling sessions, thereby enriching the narrative experience and fostering children’s creativity through facilitating multisensory feedback and emotional behaviors.
@@
These examples illuminate the multifaceted roles that robots can play in creative processes, from facilitating creative processes as tools and partners to acting as resources (Chung et al., 2021). To fully grasp the potential of robots in the creative domain, it is essential to explore how their material features and interactive capabilities shape and support creative expression. Beyond the roles that robots play in creative processes, we are also interested in what is unique about robots (e.g., material features and interaction processes) that accounts for their continual advancement and application in creative contexts in a more integrated way. The artistic appropriation of robots may provide part of the answer. Notably, robotic artists engage proactively with robots when creating artwork, exploiting the artistic potential of robots, rather than passively accommodating the robots’ inherent limitations. They break down robots, redefine them, and repurpose them as instruments for artistic goals. We propose that robots’ specific manifested qualities embed the potential to stimulate creativity and drive such interest and engagement by artists as in their practice. Thus, we aim to better explore the above questions by examining the practices and characteristics of artists’ voluntary utilization of robots in achieving artistic ends. Answering these questions can contribute to knowledge about the unique capabilities and roles of robots in creative contexts.
@@ 2.2. Robotic Art
Robotic art involves the use of robots as both mediums and subjects of artistic expression (Penny, 2016; Kac, 1997). Robotic art covers a group of practices that is usually related to computer-automated cultural artifacts (CACA) (Penny, 2013), although some authors refer to the broader field of mechanically driven kinetic art in an attempt to frame it within media art practices (Hoetzlein, 2009). From a general perspective, a collection of works with a certain level of machine autonomy is often addressed as robotic art (Kac, 2001), belonging to a broader category of machinic art, which includes ‘telerobotic art’ (Reichle et al., 2009), ‘postbiological life forms’ (Reichle et al., 2009), ‘robotic installation art’ (Shi, 2023), and so on. Artworks featuring robots not only create more engaging experiences but also challenge the boundaries of intelligent agency and human creativity. For instance, robotic artists prompt us to question whether a robot that creates art can be considered an artist in its own right (Shi, 2023; Mikalonytundefined and Kneer, 2022).
Robots have been adopted in artistic expression and creation due to their anthropomorphic qualities, creating a metaphor as social components and providing a degree of sociality (Jeon, 2017; Kahn et al., 2014). In this context, artists can leverage robots’ social behaviors to stimulate artistic expression. For instance, natural language dialogue can establish social connections with creative producers and guide creative imagination through questioning (Kahn et al., 2016; Hu et al., 2021). Artists have utilized robots’ facial expressions, body postures, movements, and speech intonation to provide emotional communication during creative theater performances, thereby encouraging the generation of creative human-human/robot interaction (Knight, 2011; Li et al., 2023). These characteristics distinguish robots from traditional artistic materials such as plaster or fabric, allowing artists to create artwork and express themselves more dynamically through interactions with living materials.
@@
Thanks to robots’ interactivity, robotic artworks often incorporate audiences into the creation process, allowing artists to harness more collective creativity and enabling greater possibilities in artistic expression (Guljajeva and Canet Sola, 2022; Mikalauskas et al., 2018). Examples include Tresset’s performance installations with Paul, a custom robot art and computer program for real-time portrait drawing (Tresset and Deussen, 2014). In these works, audiences not only observe a robot creating art but are invited to witness the creative process of a technological ensemble at work, observing how human artists and robotic systems continuously influence each other (Chung, 2022; Gomez Cubero et al., 2021; Tresset and Deussen, 2014).
While these researches have demonstrated the potential of robots in artistic creation and their capacity to stimulate creativity, they discuss discretely how creativity is supported and unleashed by single robotic systems, revealing limited approaches toward creativity. There remains a need for research that explores multiple cases involving diverse forms of robots to deepen our understanding of how robots can contribute meaningfully to the creative process. Our study addresses this gap by analyzing the creative practices of multiple robotic art practitioners working with various mediums, such as drawing, dance, and installation. This approach allows us to uncover shared experiences and insights across different artistic forms, potentially revealing broader principles for how robots can effectively facilitate and shape creative practices.
@@ 2.3. Artistic Practices in HCI
The relationship between art and technology is becoming increasingly intertwined, with artistic practices being extensively integrated into the HCI community to study human interactions with machines in specific contexts (Jeon et al., 2019). By adopting exploratory approaches, these practices challenge traditional usability-study research design. For instance, some studies involve participants in spatial interactive sound installations and use sound theory and micro-phenomenological interviews to analyze the interplay between sensory perception, bodily intent, and sound (Robson et al., 2024; Frid et al., 2019) capturing the subtleties of how users perceive and engage with sound in immersive environments that are often overlooked by conventional measurement techniques. Rajcic and McCormack (2020) introduced an emotion-responsive mirror that generates poetry in real-time, exploring the contrasts in emotional cognition, processing, and expression between machines and users. Similarly, Being the Machine by Devendorf and Ryokai (2015) invites artists to perform 3D printing manually in a machine-like manner, probing questions of agency and control in human-machine interactions. Furthermore, collaborations with choreographers have led to the creation of movement-sonification scarves as probes in dance performances, investigating dancers’ engagement with music and aiding HCI researchers in exploring how to design body-centric interaction technologies (Bang et al., 2023). The creation of artistic works, with their inherent ambiguity and potential for unforeseen human interactions, parallels the challenges in studying human interactions with emerging technologies (Coessens et al., 2009). Interactive artworks empower HCI researchers to expand their understanding of emergent human behaviors in relation to machines.
@@
However, the role of artistic practice in HCI research has been a subject of ongoing debate with some researchers questioning integrating artistic knowledge into HCI (Frankjaer and Dalsgaard, 2020; Candy, 2011). Our review suggests that examining the relationship between art and technology serves as both inspiration and a means to understand the interaction and the broader social world (Taylor et al., 2017), leading to novel design research methods such as cultural probes (Gorichanaz, 2020; Gaver et al., 1999), participatory/co-design (Clarke et al., 2016; DiSalvo et al., 2007; Soden et al., 2020), and arts-based ethnography (Kang et al., 2022). Some studies incorporate interdisciplinary teams in the creation and performance of artworks to explore nonlinear collaboration across disciplines, highlighting the intersection of technological and aesthetic thinking (Kang et al., 2022, 2018). Research on self-situated performance practices expands our understanding of new relationships and interactions between researchers and participants in HCI, allowing for ongoing refinement of research practices (Taylor et al., 2017). In these cases, artistic practice is not merely an object for investigation but something that implies broader issues of research, offering art creators, participants, and researchers opportunities for reflection and knowledge transformation.
@@
Besides, artistic practices also bring diverse aesthetic experiences to the HCI community, helping researchers focus more on the sensory dimensions of interactions with technology. The concept of aesthetic computing, exploring how art and aesthetics are applied in different areas of computer science, introduces a sensuous aesthetic framework to the exploration of technology in HCI research (Fishwick, 2006; Jeon, 2017). Artistic practices bring embodiment and interaction to the forefront, placing human sensory experience, judgment, and contemplation at the core of aesthetic experiences (Taylor et al., 2017). For example, soma design focuses on elevating the sensory experiences of bodily activities to consciousness, utilizing these experiences as resources in the design process. Designers engage deeply with materials—whether physical, digital, cultural, or movement-based—through all their senses, cultivating their aesthetic sensibilities to create innovative designs (Hook, 2018; Höök et al., 2016; Luft et al., 2023). Consequently, various artistic practices, including choreography (Allen d’Ávila Silveira et al., 2022; Eriksson et al., 2019), sound art (Kilic Afsar et al., 2023), and media art (Branch et al., 2021), are increasingly incorporated in HCI to shift the focus to bodily sensations, prompting a reevaluation of our judgments about technology.
@@
Integrating artistic practices into HCI research has shown significant potential in broadening our understanding of user experience and technology design. This study explores how robotic art can help the HCI community deepen its understanding of the relationship between human creative behaviors and robots as facilitative materials. We aim to bridge the gap between artistic practice and technology design for creativity by proposing a framework explicating how computing systems like robots serve as the tools and sites for creative endeavors and offering subsequent design implications for creative activities.